OpenStudy (anonymous):

A florist stocks red roses and white roses. Of these, some have thorns and some do not. Let R be the event that a rose is red, and let T be the event that a rose has thorns. Suppose that the following data has been collected about the roses: 1/4 of the roses have thorns, 1/5 of the red roses do not have thorns, and 3/7 of roses with thorns are white. Consider the experiment of selecting a rose at random. Use the data to find the probability of each of the four events R ∩ T R ∩ ~T ~R ∩ T ~R ∩ ~T

OpenStudy (anonymous):

For the first one I reasoned that you have a 1/4 chance of selecting a rose with thorns and of that 1/4 3/7 of them would be white. The rest of the 4/7 thorned roses must then be red. So I think P(R ∩ T ) = (1/4)*(4/7) = 1/7

OpenStudy (anonymous):

Please let me know if you think I did the first one correctly. I'm stuck on the second one

OpenStudy (anonymous):

I think for the next one you would have a 1/7 chance of picking a red rose with thorns and a 4/5 of red roses have thorns then you divide this chance by 4 to get the 1/5 chance there are no thorns P(R ∩ ~T ) = (1/7)*(1/4) = 1/28

OpenStudy (anonymous):

I think for the next one you know you have 1/4 chance of getting a rose with thorns and also that 3/7 of them are white so P(~R ∩ T ) = (1/4)*(3/7) = 3/28

OpenStudy (anonymous):

first one is right

OpenStudy (anonymous):

ok that's a good sign

hartnn (hartnn):

1/5 of the red roses do not have thorns,, doesn't that mean that R ∩ ~T=1/5 ?

OpenStudy (anonymous):

it is late and my head is spinning, but with careful thought we can figure out what proportion of the roses are red and anything else you need

OpenStudy (anonymous):

even if 1/5 of the red roses don't have thorns don't we need to consider the probability of picking a white rose as well?

OpenStudy (anonymous):

I appreciate it satellite :)

OpenStudy (anonymous):

hold on one sec, i will see if i can make it clear

OpenStudy (anonymous):

For the last one I thought that if you have a 1/4 chance of picking a rose with thorns then 3/4 must not have thorns. Of this 3/4 you have a 1/28 chance of picking a red rose from the second part. This would give you P(~R ∩ ~T) = (3/4)*(27/28) = 81/112

OpenStudy (anonymous):

I don't feel very confident about these

OpenStudy (anonymous):

ok my brain is shot, but lets try this, and then we can do it correctly with probabilities suppose there are 700 roses, then one fourth of them or 175 have thorns and the rest do not of those 175, 3/7 are white so there are 75 white roses with thorns, and therefore 100 red roses with thorns that is 4/5 of the red roses, so there must be 125 red roses all together of which 100 have thorns and 25 to not

hartnn (hartnn):

i was doing exactly that ^^

hartnn (hartnn):

so 2n will be 25/700 = 1/28, so thats correct

OpenStudy (anonymous):

so we now now the exact composition of these 700 roses 125 red and 575 white 100 red roses have thorn, 25 do not, 75 white roses have thorns, so 500 do not now you can answer any question

OpenStudy (anonymous):

Yes this is perfect. I think I confirmed all my answers except the last one as well

OpenStudy (anonymous):

oh sorry i meant \[P(R\cap T^c)=\frac{1}{7}\]

OpenStudy (anonymous):

you can redo this without the 700 and see that the answers are the same

OpenStudy (anonymous):

wouldn't it be 25/700 or 1/28? for P(R∩~T)

OpenStudy (anonymous):

yeah i told you my brain was shot

hartnn (hartnn):

yup, that i already mentioned

hartnn (hartnn):

last one comes out to be 500/700 = 5/7... ?

OpenStudy (anonymous):

last one would be \(\frac{5}{7}\) i think white without thorns try repeating the same computation without the 700

OpenStudy (anonymous):

over and out

OpenStudy (anonymous):

thanks again satellite! :D

hartnn (hartnn):

not red and thorns = white and thorns = 75/175= 3/7

hartnn (hartnn):

1/7,1/28,3/7,5/7

OpenStudy (anonymous):

wouldn't it be 75/700 or 3/28?

hartnn (hartnn):

but there are only 175 flowers with thorns....

hartnn (hartnn):

oh, wait, its 75/700,sorry

OpenStudy (anonymous):

but you need the probability of choosing from all 700. You aren't guaranteed to get one with thorns

OpenStudy (anonymous):

yeah

hartnn (hartnn):

i also get confused with this

OpenStudy (anonymous):

I just don't know how to reason my way through the last one to get 5/7

OpenStudy (anonymous):

you have a 3/4 chance to get a rose without thorns

OpenStudy (anonymous):

how do you turn that into 5/7?

OpenStudy (anonymous):

ah I got it. It's 3/4 chance minus the 1/28 chance you get a red one. So (3/4)-(1/28) = 5/7

hartnn (hartnn):

or 2/7 with thorns are red, so all others are 1-2/7 = 5/7

OpenStudy (anonymous):

yeah that works too

OpenStudy (anonymous):

thanks for your help as well hartnn!

hartnn (hartnn):

even i learned this....its been some time i haven't done such problems....

hartnn (hartnn):

all thanks to satellite

OpenStudy (anonymous):

It's my first time taking statistics, so this is all new to me haha

OpenStudy (anonymous):

I still appreciate you taking you time to take a look at this problem anyways

hartnn (hartnn):

no problem ^_^